OpenStudy (anonymous):

Which rule will find the nth term of the arithmetic sequence -58, -65, -72, -86, …?

OpenStudy (anonymous):

@CrazyMexican816

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Do you know how to do this problem ?

OpenStudy (anonymous):

yeah it might take me a few minutes I haven't done a prob like this in a while...

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

Alright, thanks so much

OpenStudy (anonymous):

Equation for arithmetic sequence a{n} = a{1} + d(n - 1) Plug in known values 2 = a{1} + -4(10 - 1) 2 = a{1} - 36 38 = a{1} a{1} = 38 Plug d = -4 and a{1} = 38 into arithmetic sequence equation a{n} = 38 - 4(n - 1)

OpenStudy (anonymous):

Hold on let me get the answer choices

OpenStudy (anonymous):

... I was kinda confused lol hope i helped

OpenStudy (anonymous):

an=-65-7n an=-65+7n an=51-7n an=-51+7n an=-51-7n

OpenStudy (anonymous):

I would try the second to last one

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

very confused with this one sorry...

OpenStudy (anonymous):

It's okay. thanks for helping!!

OpenStudy (anonymous):

no prob!